You deposit $1000 today, how much you will get in five years if Bank ABC is offering 10%

monthly compounded on its saving accounts?
Answer #1

Correct Answer:

Amount after 5 years

$ 30,44,816.40

Working:

Principal

$   10,000.00

Rate (monthly)

                 0.10

time

                       5 Years

Amount = Principal * (1 + Rate/) ^ 5*12

= 10,000 * (1 + 0.10) ^ 60

= 10,000 * (1.10) ^60

= 10,000 * (1.10) ^60

= 10,000 * 304.4816

= $ 30,44,816.40

Note: Since the interest is compounded monthly therefore time (Years) will be multiplied by 12

End of answer.

Please give a thumbs-up, it will be highly appreciated.

Thanks.